OpenStudy (anonymous):

20 - 4/5 x greater than or equal to 16 ?

OpenStudy (saifoo.khan):

\[20 - \frac45s \ge 16\]

OpenStudy (anonymous):

exactly lol

OpenStudy (anonymous):

sorry I guess I dont know how to make some signs on the keyboard

OpenStudy (saifoo.khan):

\[-\frac45s \ge 16-20\]\[-\frac45s \ge -4\]

OpenStudy (saifoo.khan):

\[\frac45s \le 4\]Cross multiply. \[4s \le 4*5\]

OpenStudy (saifoo.khan):

\[s \le 4\]

OpenStudy (anonymous):

are you sure on this answer? it gives both greater than/less than 5 and both greater or less than 32 1/2 as possible answers